Please help me with this
What is the equation for the inverse of the function f (x) = 4x - 3?
: Please help me with this
What is the equation for the inverse of the function f (x) = 4x - 3?

Answer by Fombitz(1756)

To find the inverse of a function, replace f(x) with y.
y=4x-3
The next step is interchange y for x and vice versa.
highlight(x)=4*highlight(y)-3
Then solve for y.
x=4y-3
4y=x+3
y=(x+3)/4
The inverse of f(x),
f^(-1)(x)=(x+3)/4